The density of a gas of relative molecular mass 28 at a certain temperature is 0.90 kgm-3
.
The root mean square speed of the gas molecules at that temperature is 602 ms-1
. Assuming that the rate of a gas is inversely proportional to the square root of it's density calculate the density of the gas standard temperature and pressure of it's root mean square speed is 490m/s.
1
Expert's answer
2020-07-02T17:13:44-0400

According to Graham's law, the ratio of root mean square velocities of gases at standard temperature and pressure are inversely proportional to their molecules' masses:


"\frac{v_1}{v_2}=\sqrt{\frac{m_2}{m_1}}.\\\space\\
m=\Mu n=\rho V=\rho V_mn."


Hence:


"\frac{v_1}{v_2}=\sqrt{\frac{\rho_2 n_2}{\rho_1 n_1}}."

For the same amount of substance "n_1=n_2":


"\frac{v_1}{v_2}=\sqrt{\frac{\rho_2 }{\rho_1}},\\\space\\
\rho_2=\rho_1\bigg(\frac{v_1}{v_2}\bigg)^2=0.9\bigg(\frac{602}{490}\bigg)^2=1.35\text{ kg/m}^3."